Show that the function f(x)=1/(x^2+π^2 ) can be taken as a probability density (distibution) function of a random variable X. Find p(X>π). Find also the cumulative distribution function F(x) of the random variable X. Find, finally, mean and standard deviation of the random variable X1 Show that the function f(x) = can be taken as a probability density (distibution) x²+x² function of a random variable X. Fi
Answer #1

a As f(x) 7,0 for all x 7,0 and I fulde = x da xta2 20 / dx x²+22 - 2 - o) - Thus fex) = valid RDA. canca x2 P(xya) = f fromgo T * = doesnt site. Flala tan tan Flal (tan) + 5) x dia Mean; 4= E(X) - jafx) dx = x²+72 - ll=0; is Odd function o Also E